UT 20131115 LUCI2 Nighttime Commissioning

Observer: WSeifert, JHeidt, PBuschkamp, APramskiy, KPolsterer, MEdwards, DThompson
Telescope Operator: Sallanson
LBTO Support: TSargent

Plan for the night

Weather seems OK for the first part of the night, but will become worse later on

  • Test LS-acq
  • Tests MOS-acq
  • Preset bino-mode (check with telescope)
  • Ghosts, scattered light

Summary

Did successfully a binocular preset + offset incl. taking images.

Tests for LS- and MOS-acq. LS-acq translation only works (but not translation and rotation).

Took a LS-spectrum to determine efficiency of G200.

Some closed-dome tests.

Lost the last 3hrs of the night.

Preparations

Darks
  • Dark in dome, only blind filters in, no mask in FPU
  • 0011-0035 1x2.74s LIR
  • 0036-0060 1x1.37s MER (accidentally did MER2)

Note: When running by hand Nexp > 1 it is not possible to abort the full sequence, just the current DIT.
Still, all exposures were saved. Kai said they are interrupted, read out and saved.
But it would still be better to not start the next exposure after an abort.

Ingo modifed his LS-acq tool based on our experience last night,
not clear if it works due since the problem might be hardcoded in ALADIN

Details

Files are luci2.20131114.iiii

00:24 Sunset

00:35 Open dome, Weather: Temp + 2deg, Hum = 40%, low wind
mostly clear, but quite a few Cirrus around

00:48 z-flats, 10000-cts, files #62-101

01:05 Walter aligned field-stop

01:10 We are collimated, seeing is poor, FWHM ~ 2.0", bad weather is approaching

Tests LS acquisition

01:11 Preset BS9188, PA=0, LS-acq, looks better, offset in alpha is OK now, except that the sign is wrong

01:49 Redo, manually using star and "fake" LS, works

02:01 Check for full circle, incl. star and LS in FPU, RPC error, power cycling of AGw

02:14 Resend preset, full Acq. and alignment incl LS, => OK

02:20 BS9188, PA=60, star and "fake", Ok, redo full circle => OK

02:30 BS9188, PA=231, two stars (alignment incl. rotation), rotates in the wrong direction
WCS-information in header seems to be wrong

Tool needs to be optimized: in case slit is not perfectly aligned, centering on slits
should be done on two positions to correct for tilt

03:25 Alexander found the bug in the SW, new SW-version installed, still not OK

04:15 After several revisions OK in the end

04:20 Recheck on BS9188 PA=0 if LS-acq. is still OK, yes

04:27 Recheck on BS9188 PA=231 LS-acq., OK

04:32 PA=231, two stars (alignment incl. rotation), don't work in rotation, need to change sign, did so

04:50 PA=231, alignment of stars along fake slit (incl. rotation) again => worked

05:05 PA=60, alignment of stars onto 1" slit, crazy, orientation is correct, but translations not

05:50 Repeat at PA=231, works....but only for star close to center of rotation
=> Tool uses first star (in case of translation+rotation) as center of rotation
(should us crpix-values from header)

Tests MOS acquisition

06:14 Lets try MOS-acq on UDS3H1, doesn't work, Ingo needs to check

Binocular preset + offset, taking images

06:59 Preparation for bino-preset by Steve

07:22 Binocular preset, BS9107, including offsets on source, SUCCESSFULL

07:35 Another bino-preset by Alexander, but failed, needs to be analyzed

Efficiency G200

07:57 Check for efficiency: Preset BS9107 in spectro-mode, script Persson_BS9107
Note: For the moment being, after a bino-preset, the telescope service/manager/scheduler
on LUCI-side needs to be restartet

08: 10 Start taking spectroscopic data on BS9107, works, slitview #195, HK_spec@1.93mue, 2*(3*30sec int), files #196/197
zJ_spec@1.1mue, 2*(3*30sec, erroneously saved individually), files #198-257, 2*(3*30sec int), files #258,259,
hK_spec@1.93mue, 2*(3*30sec), files; alignment and slitview with 3.75 camera, spectra with 1.8camera

*ALADIN error message: warning: two different pixel scale values detected

08:40 Preset HIP17971, not possible going into the wind (wind close to the limit)

Lots of cirrus now out there, extinction about 1mag

09:28 Telluric HIP17718, LS 2.0", acq. #265, slitview #266, slitview after alignment, #267
HK_spec, 1.91mue, zJ_spec, 1.1mue, 2*(2*30sec int each), #268-271

10:00 Dave is performing a "Spin"-test to check on whether the AGw is in which condition

Ghost tests

10:13 Michelle pointed to a bright M5-star to check for ghosts, error on FW2 when moving to Ks, fixed by Kai

10:17 Did some ghost tests, files #272-288

10:46 Closed dome due to high humidity

Closed-dome activities

11:31 Camera-repro, Rot=210, files #289-298
11:39 Camera-repro, Rot=300, files #299-308
11:49 Camera-repro, Rot=120, files #309-318
12:00 Camera-repro, Rot=30, files #319-328

12:17 Grat_select_repro, Rot=210, files #329-336, got stuck, restart scheduler
12:34 Grat_select_repro, Rot=210, files #337, FW2 got stuck again, stopped.

13:44 End of nighttime activities Com#1 LUCI2.

-- JochenHeidt - 15 Nov 2013
Topic revision: r9 - 05 Dec 2013, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback